what is the area of a triangle if the base is 16 m and the height is 3m?
Answer by nerdybill(7384)

what is the area of a triangle if the base is 16 m and the height is 3m?
.
area of any triangle is (1/2)bh
where
b is base
h is height
.
area = (1/2)bh
area = (1/2)(16)(3)
area = (8)(3)
area = 24 square meters